# Mixed-Strategy Nash Equilibrium for Crowd Navigation

### v2

@@ II-D Game-theoretic planning
Even though game-theoretic planning falls into the category of coupled prediction and planning, here we separately discuss this body of work given its close connection to our work. The key difference between game-theoretic methods and other coupled-prediction-and-planning methods is that game theory assumes each agent makes individual optimal decisions and it provides a rigorous optimality criterion for decision-making—the concept of equilibrium. The two most commonly used equilibrium notions are Nash equilibrium [10] and Stackelberg equilibrium [58]. Stackelberg equilibrium, in general, has lower computation complexity since it enforces a leader-follower structure. On the other hand, Nash equilibrium is considered as the more rigorous and natural notion, since it assumes equal status for all agents and all agents make decisions simultaneously.
@@
In order to apply game theory models, crowd navigation is often formulated as a dynamic game: agents, with states governed by certain dynamics, optimize control policies as pure-strategies. One of the pioneering works in this direction is [9], where the interaction between a human-driven vehicle and an autonomous vehicle is modeled as a dynamic game. Since then, dynamic game solvers with better numerical efficiency have been developed. [59] formulates a general-sum dynamic game for crowd navigation, where the local Nash equilibrium is solved through the combination of an iterative best response scheme and iterative linear-quadratic regulator. [60] proposes a fast augmented Lagrangian solver that converges locally to the generalized Nash equilibrium of a dynamic game, which supports real-time model predictive control for autonomous driving. However, pure-strategy Nash equilibrium expects humans to react exactly as predicted, which is too assertive for interacting with humans. Furthermore, existing pure-strategy game solvers can only approximate a locally optimal solution, since computing the Nash equilibrium of a generalized game is computationally intractable [12].
@@
To address the lack of flexibility in pure-strategy Nash equilibrium, probabilistic variants of Nash equilibrium have been proposed. As a pioneer work in this direction, [61] proposes a stochastic dynamic game formula that lifts the strict rationality assumption of Nash equilibrium and instead assumes bounded rationality of agents. The authors propose a new notion of equilibrium named Entropic Cost Equilibrium (ECE) and show that ECE is equivalent to the mixed-strategy Nash equilibrium of a maximum entropy game. The proposed equilibrium formula also enables inverse inference of interaction policy from observations. In [62, 63], the multi-modality in stochastic games is investigated from the perspective of linear-quadratic games and partially observable Markov decision process (POMDP). In [64], an explicit mixed-strategy game formula is proposed, allowing agents to simultaneously optimize for multi-agent pure-strategies, but this formula does not recover the mixed-strategy Nash equilibrium.
@@
Lastly, game theory models can be combined with other methods for online adaptation during crowd navigation. For example, Nash equilibrium can be used to infer other agents’ internal states, such as altruism. [65] proposes to use Nash equilibrium as an inference model for online estimation of human drivers’ social value orientation, leading to the improved prediction accuracy of human driver actions during the interaction. Similar ideas are also explored in [66, 67, 3]. Another example of bootstrapping adaptive decision-making with game-theoretic models is [68], where the game-theoretic model is used for providing candidates for contingency planning.
Limitations of existing game-theoretic planning methods
@@
There are two major limitations of existing game-theoretic planning methods for crowd navigation. The first limitation is the high computation complexity, which prevents existing methods from being applied to real-time navigation in human crowds. For example, as one of the fastest pure-strategy dynamic game solvers, ALGAMES [60] still suffers from a time complexity of $O(M^{3})$ with $M$ being the number of players, thus can only perform real-time inference with no more than 2 human agents. The computation of mixed-strategy Nash equilibrium is even more burdensome. In [61], the method is also only demonstrated with no more than 2 human agents. In [64], the computation of mixed-strategies is offloaded to a neural network training offline and this approach is only demonstrated for two-agent interaction. The second limitation is the lack of flexibility in representing uncertainty in strategies. In [61] is based on the Linear-Quadratic-Gaussian formula. In [64], the mixed-strategy model is limited to a small finite number of pure-strategies, instead of a probability distribution.
@@
We argue that these two limitations are largely due to the “top-down” framework taken by existing methods. In the “top-down” framework, multi-agent interaction is formulated and solved as a generalized game, where no assumption of agent objective is made. The agent objective will be specified later for specific applications such as crowd navigation. While this “top-down” approach could be applied for arbitrary types of multi-agent interaction beyond crowd navigation, it suffers from high computation complexity and often requires narrowing down strategy representation (e.g., Gaussian distributions). Our work, however, takes a “bottom-up” approach—we tailor a stochastic game formula specifically for crowd navigation, based on unique properties in crowd navigation, such as the assumption that each agent has an individual goal-oriented objective while sharing a collision avoidance objective with other agents. This “bottom-up” approach allows us to apply calculus of variations [69] to compute mixed-strategy Nash equilibrium in closed form, which further enables real-time inference with a larger number of agents (up to 8 agents on a laptop CPU). In addition, our algorithm and proofs support arbitrary types of distribution and collision avoidance cost functions, which enriches the representation flexibility of our model.
III Bayes’ rule for crowd navigation
In this section, we introduce a simple Bayesian updating scheme for crowd navigation.
@@ III-A Notations and definitions
We assume there are $N$ agents, including the robot and the pedestrians, in a two-dimensional navigation environment. We start by defining a set of unique indices =$\mathcal{I}=\{1,2,\dots,N\}$ for all agents, where index $1$ is reserved for the robot. The state space of each agent is denoted as $\mathcal{X}\subset\mathcal{R}^{2}$, as we are primarily interested in agents’ planar positions. We model agent $i$’s pure-strategy as a continuous trajectory $s_{i}:\mathbb{R}^{+}\mapsto\mathcal{X}$, which maps time to the agent’s state at that moment. The space of all possible pure-strategies is called trajectory space, denoted as $\mathcal{S}$.
Remark 1.
We do not address the perception aspect of the crowd navigation problem. In this work, we assume the robot has reasonably accurate estimations of its pose and nearby pedestrians’ positions.
@@
(Mixed-strategy) We model a mixed-strategy as a probabilistic belief over an agent’s possible pure-strategies. We denote agent $i$’s mixed-strategy as:
$p_{i}(s):\mathcal{S}\mapsto\mathbb{R}_{0}^{+}.$ (1)
Lemma 1.
Since a mixed-strategy is a probability distribution, it must satisfy the following property:
@@
(Nominal mixed-strategy) We define each agent has a nominal mixed-strategy before interacting with other agents, denoted as $p^{\prime}_{i}(s)$.
@@
We assume the nominal mixed-strategy represents the agent’s intent without the presence of other agents, thus it does not reflect collision avoidance behavior.
@@
(Collision risk) Collision risk is defined as a function $r(s_{1},s_{2}):\mathcal{S}\times\mathcal{S}\mapsto\mathbb{R}_{0}^{+}$, which evaluates the collision risk between two pure-strategies (trajectories).
Assumption 1.
We assume the collision risk function’s support is bounded, close, and compact.
@@
(Expected collision risk) Given one agent’s mixed-strategy $p_{i}$, the expected risk of another given pure-strategy $s$ colliding with the agent is:
@@
(Joint expected collision risk) The joint expected collision risk between two agents is defined as the joint expectation of collision risk with respect to their mixed-strategies:
@@
Figure 3: Examples of the negotiation convergence process during our iterative Bayesian updating scheme. (a) Two-agent negotiation in one-dimensional space. (b) Two-agent hallway passing. Here the mixed-strategy is modeled in the space of trajectories. (c) Four-agent crossing.
III-B Two-agent social navigation
We start with describing the iterative Bayesian updating scheme that finds the mixed-strategy Nash equilibrium between two social navigation agents. Bayesian updating requires two key components: a prior belief and a conditional likelihood. Here the prior belief is each agent’s current mixed-strategy, denoted as $p_{i}^{\prime}(s)$. By assuming each agent follows the expected utility hypothesis [70], we propose the following conditional likelihood function to reflect collision avoidance behavior.
@@ III-B Two-agent social navigation
(Conditional likelihood) Given the mixed-strategy of agent $j$, the likelihood of measuring agent $i$ following a pure-strategy $s$ is defined as:
$z(s|p_{i})$ $=\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}.$ (5)
@@
The use of inverse exponential in (5) is an empirical design choice, inspired by Gaussian distributions. But as will be shown later, the inverse exponential is necessary for the convergence to Nash equilibrium.
@@
(Bayesian posterior strategy) Given the mixed-strategy $p_{i}^{\prime}(s)$ of agent $i$ and the mixed-strategy $p_{j}(s)$ of agent $j$, the Bayesian posterior mixed-strategy of agent $i$ after applying Bayes’ rule is:
@@
$=\eta\cdot p_{i}^{\prime}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}$ (7)
@@
We now describe the iterative Bayesian updating algorithm for two-agent social navigation. Figure 3(a) and Figure 3(b) show iterations of two-agents’ mixed-strategies in an illustrative one-dimensional space and a hallway passing scenario, respectively.
Algorithm 1 Two-agent social navigation
1:procedure TwoAgentSocialNav($p_{i}^{\prime},p_{j}^{\prime}$)
@@
Algorithm 2 Multi-agent crowd navigation
1:procedure MultiAgentSocialNav($p^{\prime}_{1},\dots,p^{\prime}_{N}$)
@@
8:              $p^{[k]}_{/i}\leftarrow\frac{1}{N-1}\left[{\sum_{j<i}}p^{[k+1]}_{j}{+}\sum_{j>i}p^{[k]}_{j}\right]$
9:              $p_{i}^{[k+1]}\leftarrow\eta\cdot p^{\prime}_{i}\cdot z\left(s|p^{[k]}_{/i}\right)$
@@
III-C Multi-agent crowd navigation
@@
(Conditional likelihood (multi-agent)) Given the mixed-strategies of all agents other than $i$, the likelihood of measuring agent $i$ following a pure-strategy $s$ is defined as:
$z(s|p_{/i})$ $=\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}.$ (8)
$p_{/i}(s)$ $=\overline{\sum_{j\in\mathcal{I}/i}}p_{j}(s)$ (9)
where $\overline{\sum}$ denotes averaged sum.
@@
(Bayesian posterior strategy (multi-agent)) Given the mixed-strategy $p_{i}^{\prime}(s)$ of agent $i$ and the mixed-strategies of all the rest of the agents, the Bayesian posterior mixed-strategy of agent $i$ after applying Bayes’ rule is:
@@
$=\eta\cdot p_{i}^{\prime}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}$ (11)
@@
We now describe the iterative Bayesian updating algorithm for multi-agent crowd navigation in Algorithm 2. Figure 3(c) shows iterations of multi-agent mixed-strategies in a crossing scenario.
@@
Across all the examples in Figure 3, we want to point out two important properties of our algorithm: (1) Bayesian updating allows non-symmetric and multi-modal mixed-strategies during iterations, which expands the expressiveness of our method; (2) Although the iterative Bayesian updating scheme does not simultaneously update all agents’ strategies, the converged mixed-strategies can still be symmetrical. As we will discuss in the next section, the symmetry comes from the fact that our algorithm converges to a global Nash equilibrium. A detailed demonstration of mixed-strategy evolution during the iterative Bayesian updating scheme can be found in Extension 1.
IV Mixed-strategy Nash equilibrium for social navigation

### v4

@@ II-D Game-theoretic planning
Even though game-theoretic planning falls into the category of coupled prediction and planning, we separately discuss this body of work here, given its close connection to our work. The key difference between game-theoretic methods and other coupled prediction and planning methods is that game theory assumes each agent makes individual optimal decisions and provides a rigorous optimality criterion for decision-making—the concept of equilibrium. The most commonly used equilibrium notions are Nash equilibrium [10] and Stackelberg equilibrium [58]. Stackelberg equilibrium, in general, has lower computation complexity since it enforces a leader-follower structure. On the other hand, Nash equilibrium is considered the more rigorous and natural notion for crowd navigation, since it assumes equal status for all agents and all agents make decisions simultaneously.
@@
To apply game theory models, crowd navigation is often formulated as a dynamic game: agents, with states governed by certain dynamics, optimize control policies as pure strategies. One of the pioneering works in this direction is [9], where the interaction between a human-driven vehicle and an autonomous vehicle is modeled as a dynamic game. Since then, dynamic game solvers with better numerical efficiency have been developed. In [59], the authors formulate a general-sum dynamic game for crowd navigation, where the local Nash equilibrium is solved by combining an iterative best response scheme and linear-quadratic regulator. In [60], a fast augmented Lagrangian solver is proposed, which converges locally to the generalized Nash equilibrium of a dynamic game and supports real-time model predictive control for autonomous driving. However, pure strategy Nash equilibrium models expect humans to react exactly as predicted, which is too assertive for interacting with humans. Furthermore, existing pure strategy game-theoretic planners can only approximate a locally optimal solution since computing the Nash equilibrium of a generalized game is computationally intractable [12].
@@
Probabilistic variants of pure strategy Nash equilibrium have been proposed to address the lack of flexibility. In [61], an iterative best response scheme is combined with a model predictive path integral (MPPI) control framework to approximate cooperative stochastic control policies between two autonomous vehicles. However, this approach does not formally guarantee recovering a mixed strategy Nash equilibrium. In [62], a stochastic dynamic game formula is proposed to lift the strict rationality assumption of Nash equilibrium and instead assumes bounded rationality of agents. The authors propose a new notion of equilibrium named Entropic Cost Equilibrium (ECE) and show that ECE is equivalent to the mixed strategy Nash equilibrium of a maximum entropy game. The proposed equilibrium formula also enables inverse inference of interaction policy from observations. In [63] and [64], the multimodality in stochastic games is investigated from the perspective of linear-quadratic games and partially observable Markov decision process (POMDP), where the uncertainty originated from the partially observable objectives of non-ego agents. In [65], an explicit mixed strategy game formula is proposed, allowing agents to simultaneously optimize multiple multi-agent pure strategies, but this formula does not recover the mixed strategy Nash equilibrium.
@@
Lastly, game theory models can be combined with other methods for online adaptation during crowd navigation. For example, Nash equilibrium can be used to infer other agents’ internal states, such as altruism. In [66] the pure strategy Nash equilibrium model is used as an inference model for online estimation of human drivers’ social value orientation, leading to the improved prediction accuracy of human driver actions during the interaction. Similar ideas are also explored in [67], [68], and [3]. Another example of bootstrapping adaptive decision-making with game-theoretic models is [69], where the game-theoretic model provides candidates for contingency planning.
Limitations of game-theoretic planning methods
@@
There are two major limitations of existing game-theoretic planning methods for crowd navigation. The first limitation is the high computation cost, which prevents existing methods from being applied to real-time navigation in human crowds. For example, as one of the fastest pure strategy dynamic game solvers, ALGAMES [60] still suffers from a time complexity of $O(M^{3})$ with $M$ being the number of players, thus can only perform real-time inference with no more than 2 human pedestrians. The computation of mixed strategy Nash equilibrium is even more burdensome. In [62], the method is also only demonstrated with no more than 2 human agents. In [65], the computation of mixed strategies is offloaded to a neural network trained offline and this approach is only demonstrated for two-agent interaction. The second limitation is the lack of flexibility in uncertainty representation. In [62], the solution is based on the Linear-Quadratic-Gaussian formula. In [65], the mixed strategy model is limited to a small number of pure strategies, instead of a probability distribution.
@@
We argue that these two limitations are largely due to existing methods’ “top-down” approach. In the “top-down” approach, multi-agent interaction is formulated and solved as a generalized game, where no specification of agent objective is made. The agent objective will be specified later for specific applications such as crowd navigation. While this “top-down” approach could be applied for arbitrary types of multi-agent interaction beyond crowd navigation, it suffers from high computation complexity and often requires narrowing down strategy representation (e.g., Gaussian distributions). Our work, however, takes a “bottom-up” approach—we tailor a stochastic game formula specifically for crowd navigation, based on unique properties in crowd navigation, such as the insight that each agent has an individual goal-oriented objective while being coupled with other agents through a collision avoidance objective. This “bottom-up” approach allows us to compute mixed strategy Nash equilibrium in closed form, enabling real-time inference with more agents (up to 8 agents on a laptop CPU, 5 agents on a low-power embedded computer). In addition, our algorithm and proofs support arbitrary types of distribution and collision avoidance cost functions, which further enriches the representation flexibility of our model.
III Bayesian update for crowd navigation
In this section, we introduce the iterative Bayesian update scheme for crowd navigation. We will focus on the algorithm description and the model predictive navigation framework, leaving the formal properties of the algorithm, such as the guaranteed convergence to a Nash equilibrium, in Section IV. For the same reason, we leave the introduction of the complete game-theoretic formula in Section IV as well.
@@ III-A Notations and definitions
We assume there are $N$ agents, including the robot and the pedestrians, in a two-dimensional navigation environment. We start by defining a set of unique indices =$\mathcal{I}=\{1,2,\dots,N\}$ for all agents, where the first index $1$ is reserved for the robot. The state space of each agent is denoted as $\mathcal{X}\subset\mathbb{R}^{2}$, as we are primarily interested in agents’ planar positions.
While we will fully introduce the game theory formula in Section IV, we introduce the concept of strategy here. A strategy describes an agent’s decision-making process. In the context of navigation, we assume agents make decisions regarding trajectories, which outline where they intend to travel over a given time period $[0,T]$.
@@
(Pure strategy) A pure strategy $s(t)$ is defined as a trajectory, which is specified as a continuous mapping from time to a state in the navigation space:
$s(t):[0,T]\mapsto\mathcal{X}.$ (1)
Definition 2.
(Strategy space) The strategy space is defined as the space of all feasible pure strategies that an agent may consider, we denote it as $\mathcal{S}$.
Definition 3.
(Mixed strategy) A mixed strategy is a probability distribution over the strategy space $\mathcal{S}$, represented as a probability density function $p(s)$:
$p(s):\mathcal{S}\mapsto\mathbb{R}_{0}^{+}$ (2)
@@
(Mixed strategy space) The mixed strategy space, denoted as $\mathcal{P}$, is the space of all probability density functions over the strategy space $\mathcal{S}$.
Definition 5.
(Nominal mixed strategy) Each agent is assumed to have a nominal mixed strategy before interacting with other agents, denoted as $p^{\prime}_{i}(s)$, with $i$ being the agent index.
@@
We assume the nominal mixed strategy represents the agent’s intent without the presence of other agents. Thus, it does not reflect any collision avoidance behavior.
@@
(Collision risk) Collision risk is defined as a function $r(s_{1},s_{2}):\mathcal{S}\times\mathcal{S}\mapsto\mathbb{R}_{0}^{+}$, which evaluates the collision risk between two pure strategies (trajectories).
@@
(Expected collision risk) Given agent $i$’s mixed strategy $p_{i}(s)$, the expected risk of another pure strategy $s(t)$ colliding with agent $i$ is:
@@
(Joint expected collision risk) The joint expected collision risk between two agents is defined as the joint expectation of collision risk with respect to their mixed strategies:
@@
Figure 2: Examples of the iterative Bayesian update process. (a) Two-agent negotiation in one-dimensional space. (b) Two-agent hallway passing, where the mixed strategy is visualized as trajectory samples. (c) Four-agent crossing.
III-B Iterative Bayesian update with two agents
We describe the iterative Bayesian update scheme that finds cooperative mixed strategies between two navigation agents. Bayesian update requires two key components: a prior belief and a conditional likelihood function—the latter is often interpreted as a measurement model. Here, the prior belief is each agent’s nominal mixed strategy $p^{\prime}(s)$. By assuming each agent follows the expected utility hypothesis [70], we propose the following conditional likelihood function to reflect collision avoidance behavior.
@@ III-B Iterative Bayesian update with two agents
(Conditional likelihood) Given the mixed strategy of agent $j$, the likelihood of measuring another agent following a pure strategy $s(t)$ is defined as:
$z(s|p_{j})$ $\propto\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)}.$ (6)
@@
The inverse exponential in (6) is an empirical design choice, inspired by Gaussian distributions. But as will be shown later, the inverse exponential is necessary for the algorithm to converge to Nash equilibrium.
@@
(Bayesian posterior strategy) Given the nominal mixed strategy $p_{i}^{\prime}(s)$ of agent $i$ and the current mixed strategy $p_{j}(s)$ of agent $j$, the Bayesian posterior mixed strategy of agent $i$ after applying Bayes’ rule is:
@@
$=\eta\cdot p_{i}^{\prime}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{j}}[r](s)\Big{)},$ (8)
@@
With the Bayesian posterior strategy defined in (8), we now describe the complete iterative Bayesian update algorithm with two agents in Algorithm 1. Figure 2(a) and Figure 2(b) show evolution of two agents’ mixed strategies in an illustrative one-dimensional example and a hallway passing scenario, respectively.
Algorithm 1 Two-agent iterative Bayesian update
1:procedure TwoAgentUpdate($p_{i}^{\prime},p_{j}^{\prime}$)
@@
Algorithm 2 Multi-agent iterative Bayesian update
1:procedure MultiAgentUpdate($p^{\prime}_{1},\dots,p^{\prime}_{N}$)
@@
8:              $p^{[k]}_{/i}\leftarrow{\sum_{j<i}}p^{[k+1]}_{j}{+}\sum_{j>i}p^{[k]}_{j}$
9:              $p_{i}^{[k+1]}\leftarrow\eta\cdot p^{\prime}_{i}\cdot z\left(s|p^{[k]}_{/i}\right)$triangleright See (9)
@@
III-C Iterative Bayesian update with multiple agents
@@
(Conditional likelihood (multi-agent)) Given the mixed strategies of all agents other than $i$, the likelihood of measuring agent $i$ following a pure strategy $s(t)$ is defined as:
$z(s|p_{/i})$ $\propto\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)}.$ (9)
$\mathbb{E}_{p_{/i}}[r](s)$ $=\sum_{j\in\mathcal{I}/i}\mathbb{E}_{p_{j}}[r](s).$ (10)
@@
(Bayesian posterior strategy (multi-agent)) Given the nominal mixed strategy $p_{i}^{\prime}(s)$ of agent $i$ and the current mixed strategies of all the rest of the agents, the Bayesian posterior mixed strategy of agent $i$ after applying Bayes’ rule is:
@@
$=\eta\cdot p_{i}^{\prime}(s)\cdot\exp\Big{(}{-}\mathbb{E}_{p_{/i}}[r](s)\Big{)},$ (12)
@@
The complete iterative Bayesian update algorithm for multi-agent crowd navigation is described in Algorithm 2. Figure 2(c) shows the evolution of multi-agent mixed strategies in a crossing scenario. We name our algorithm Bayes’ Rule Nash Equilibrium (BRNE).
@@
Across all the examples in Figure 2, we want to highlight two important properties of our algorithm: (1) Bayesian update could capture non-symmetric and multimodal mixed strategies, which expands the expressiveness of our method; (2) Although the iterative Bayesian update scheme does not simultaneously update all agents’ strategies in each iteration, the converged mixed strategies can still be symmetrical.
Figure 3: Illustration of specifying a nominal mixed strategy with a Gaussian process (GP). (a) We first specify a trajectory as the mean function of the GP. For the robot, it would be a trajectory toward the goal, generated by a meta-planner; (b) We then specify the covariance kernel parameters, either learned from datasets or hand-tuned, which give us the GP prior distribution; (c) The GP prior is insufficient as the nominal mixed strategy. It needs to be conditioned at specific time steps with user-specified marginal uncertainty. In the figure, we condition the GP prior on the first and last time step (the specified marginal uncertainty is shown as the red ellipse); the resulting distribution is the robot’s nominal mixed strategy.
Figure 4: Illustration of the model predictive crowd navigation framework. (a) The robot first takes measurements of nearby pedestrians’ positions and velocities; (b) The robot then generates the mean functions of the Gaussian processes for the pedestrians and for itself; (c) The robot specifies the Gaussian processes as the nominal mixed strategies for all agents and draws trajectory samples from them; (d) The weights of the trajectory samples are updated based on Algorithm 2 until convergence; (e) The mean of the robot’s converged mixed strategy becomes the robot’s planned trajectory.
III-D Iterative Bayesian update for model predictive crowd navigation
We now apply the iterative Bayesian update scheme in Algorithm 2 as a model predictive crowd navigation framework. The key insight is that the mixed strategy of each agent can be modeled through trajectory samples, with which Algorithm 2 is essentially a weight update scheme for the samples, allowing fast computation in practice. Below, we introduce the practical algorithmic design.
Motivation for sample-based representation
Computing the analytical mixed strategies following the Bayesian update in Algorithm 2 is intractable in practice, as the mixed strategies are represented as continuous probability density functions in the analytical formula. This is a common issue among Bayesian filter methods, as computing the analytical Bayesian posterior is, in general, intractable. Importance sampling [71] is one of the most widely used solutions to approximate the posterior from a Bayesian update efficiently. It represents the prior belief using samples drawn from the prior belief, then it computes the posterior belief by updating the sample weights based on the measurement model. With the sample weights updated, the samples can be updated further through resampling methods, such as rejection sampling [72], for the next iteration. We apply importance sampling to approximate Algorithm 2 in practice. The mixed strategies are represented as samples and weights are computed based on the Bayesian update step in Algorithm 2. Similar sampling-based representation is also used in stochastic optimal control, such as path integral control [73].
Algorithm overview
In the proposed sampling-based model predictive crowd navigation framework, the robot repeatedly (1) observes the pedestrian position and velocity, (2) generates trajectory samples to represent the nominal mixed strategies for the pedestrians, (3) generates trajectory samples toward the navigation destination to represent the nominal mixed strategy for itself, (4) applies Algorithm 2 to update the weights of the trajectory samples, and (5) computes the optimal control signal by tracking the weighted average trajectory from the robot’s converged mixed strategy. The overall process is visualized in Figure 4.
Gaussian process mixed strategy model
One of the requirements for the sampling-based Bayesian update is that we can draw samples from the prior belief, also often known as the proposal distribution, which in Algorithm 2 is the nominal mixed strategy of each agent. We model the nominal mixed strategy for each agent using Gaussian processes [74]. A Gaussian process is an infinite-dimensional normal distribution over continuous functions. In our case, these continuous functions are the trajectories, as we model a trajectory as a continuous mapping from time to the agent state. A Gaussian process is characterized by a mean function and a covariance kernel function. We use a constant velocity model as the mean function for each pedestrian, assuming the pedestrian will keep the current velocity during the robot’s planning horizon [^2] In [28], the constant velocity model is shown to be competitive against state-of-the-art learning-based models for trajectory prediction [28].. A high-level meta-planner generates the mean function for the robot as a trajectory from the robot’s current location toward the navigation goal. With the presence of static obstacles, the mean function trajectory should also avoid the obstacles. Such trajectory planning problems have been well-studied and can be solved using methods such as RRT or trajectory optimization. The parameters of the covariance kernel function can be learned from existing datasets through standard inference techniques such as maximum likelihood estimation [6] or hand-tuned if the number of parameters is small (e.g., with radial basis kernels). Once a Gaussian process is specified, following the same step in Gaussian process regression (Eq 2.24 in [74]), we can condition the specified Gaussian process on discrete time steps. This conditioning happens offline and converts the infinite-dimensional Gaussian process to a finite-dimensional normal distribution with each dimension representing a time step. It also allows us to specify the marginal uncertainty at specific time steps. The converted multivariate normal distribution is an agent’s nominal mixed strategy in practice, from which trajectory samples can be drawn efficiently during runtime, as the sampling of trajectories is now equivalent to sampling from a normal distribution. Figure 3 illustrates the steps of specifying a nominal mixed strategy with a Gaussian process. Lastly, Gaussian processes do not explicitly model agents’ dynamics. Instead, the Gaussian process kernel preserves the geometrical smoothness of the trajectory samples. Similar Gaussian process-based trajectory distribution representations have been verified in other works for crowd navigation [6] and motion planning [75].
Remark 3.
Even though we use a Gaussian process-based model for the mixed strategy, it is not the only choice. For pedestrians, the nominal mixed strategy can be the predicted trajectory samples drawn from neural network-based trajectory prediction frameworks, such as Trajectron [26] or TrajNet [76]. For the robot, instead of directly drawing samples in the trajectory space, we can also draw samples in the space of control signals by randomly perturbing a nominal control signal, similar to the widely used model predictive path integral (MPPI) control framework [73, 77]. We choose Gaussian processes for computation and sampling efficiency. Inference with Gaussian processes, which computes both the mean function and the covariance kernel from existing data, is known to be computationally expensive. However, we avoid performing Gaussian process inference by specifying the mean functions and the kernel parameters in advance. Therefore, drawing trajectory samples is equivalent to drawing samples from a multivariate normal distribution, which is affordable for real-time computation and faster than inference from neural networks.
Weight update for samples
Once the trajectory samples are generated, we can compute the conditional likelihood function (9) for each sample as the updated weight, where the continuous integral can be approximated using Monte-Carlo integration based on the samples. After normalizing the weights, the now-weighted trajectory samples represent the Bayesian posterior in (12). We repeat this process until convergence, and the weighted average trajectory from the robot’s converged mixed strategy will be the robot’s planned trajectory, from which the robot will compute control signals to track it. Note that, even though Algorithm 2 requires only one agent’s mixed strategy to be updated at a time, the weights of the trajectory samples from the same agent can be updated simultaneously, which can benefit from parallel computation for better computation efficiency. Pseudocode for an importance sampling-based implementation and a rejection sampling-based implementation are included in the appendix.
Algorithm complexity
We analyze the computational time complexity of a single iteration within the sampling-based algorithm. Given $T$ time steps as the planning horizon, $N$ agents, and $M$ samples for each agent, the complexity of computing the weights for one agent’s mixed strategy is $O(TMN)$. For each iteration of the Bayesian update scheme, the weights need to be computed for all the agents, which leads to an overall complexity of $O(TMN^{2})$ for one iteration. Note that, given $N$ agents, $O(N^{2})$ is the minimal complexity required for reasoning over all possible two-agent interaction pairs. The computation complexity of our algorithm with respect to the number of agents is lower than the state-of-the-art dynamic game solver ALGAMES [60], which has a computation complexity of $O(TN^{3})$.
IV Mixed strategy Nash equilibrium for crowd navigation